OpenStudy (anonymous):

how do I find the average percentage for 17.1,17.3,17.4,16.5

sam (.sam.):

How many numbers are there?

sam (.sam.):

\[\large \bar x=\frac{17.1+17.3+17.4+16.5}{4} \times 100\]

OpenStudy (anonymous):

@sam why *100???

sam (.sam.):

Because we are finding the percentage, so, when you want percent you multiply by 100, but if you want in terms of angle, then multiply by 360

OpenStudy (anonymous):

Ok i think I get it, thank you so much

sam (.sam.):

you sure? alright, yw

OpenStudy (unklerhaukus):

are the numbers in your question already percentages {17.1%, 17.3%,...} ?

sam (.sam.):

hmm could be that^ too

OpenStudy (unklerhaukus):

if they are already percentages, you wont need to multiply by 100,

sam (.sam.):

\[\large =\frac{17.1+17.3+17.4+16.5}{4} =17.075 percent\]

sam (.sam.):

yeah

OpenStudy (unklerhaukus):

\[\large\%\]

sam (.sam.):

\[%\]

sam (.sam.):

Percent doesn't work on latex?

OpenStudy (unklerhaukus):

``` \[\%\] ```

sam (.sam.):

\[\%\]

sam (.sam.):

alright thanks :)

OpenStudy (unklerhaukus):

the % symbol is used for comments in latex, ie stuff to the right wont appear of on rendered tex ``` \[100\% % comments\] ``` \[100\% % comments\]

sam (.sam.):

Interesting